# On the intersection of Cantor sets and products of random matrices

### v1

@@ 1.1  Background
$K_{i}=\left\{\sum_{n=0}^{\infty}\frac{d_{n}}{b^{n}}\hskip 2.0pt\middle|\hskip 2.0ptd_{n}\in D_{j}\right\}\subset\mathbb{T}=\mathbb{R}/\mathbb{Z}.$
@@
Let $\mu$ be a probability measure on $\mathbb{T}=\mathbb{R}/\mathbb{Z}$, invariant and ergodic for $T_{b}x=bx\mathrm{\hskip 5.0ptmod\hskip 3.0pt}1$. Define $2$ by $2$ matrices $A_{0},\ldots,A_{b-1}$ by
@@
This $\lambda$ is called the (top) Lyapunov exponent of the product of random matrices $A_{0},\ldots,A_{b-1}$. The RHS of (1.1) may not always exist, but it exists and is constant for $\mu$-a.e. $t$ ([2]). Also, the original statement in [8, Theorem 1.2] is formulated solely for the Lebesgue measure. However, more general statement in [8, Corollary 5.8] guarantees that this is true for any invariant and ergodic probability measure.
@@
A probability measure $\nu$ on [1,1][-1,1] is said to be $\bm{\mu}$-stationary if
$\int_{[-1,1]}g\,d\nu=\int_{[-1,1]}\left(\int_{\mathbb{T}}g\circ f_{t_{1}}(x)\,d\mu(t)\right)\,d\nu(x)$
@@
Let $\mu$ be the Lebesgue measure, and $\nu$ be any $\mu$-stationary measure. Then, $\nu$ is discrete (i.e. it is a countable sum of dirac measures), and therefore unique. The Lyapunov exponent can be calculated as
$\lambda=\frac{1}{6}\log{\frac{2}{3}}+\sum_{k=0}^{\infty}\frac{1}{4^{k+1}}\log{\frac{(3\cdot 2^{k})!}{(2^{k+1})!}}=0.797435\ldots.$
This gives the dimension to be $0.575228\ldots$.
@@
The pair $(D_{1},D_{2})$ is degenerate if and only if one of the matrices $A_{i}$ has rank one.
When this happens, the image of the corresponding $f_{i}$ is a singleton, making the $\mu$-stationary measure discrete (and thus unique). We can then calculate the dimension by tracking the trajectory of the said singleton under all $f_{i}$. The calculations by Kenyon and Peres (such as Example 1.4) fall under this category. See the recent paper by Fan and Vebitskiy [4] for a general formula under this assumption.
@@ 1.2  Main results with examples
Let $\mu=\left(0,\frac{1}{2},0,\frac{1}{2},0,0,0\right)^{\otimes\mathbb{N}}$. In this case, stationary measures are continuous, so the method by Kenyon and Peres does not work. Take any $\mu$-stationary measure $\nu$ and let $\xi_{n}=\int\phi^{n}d\nu$. Then, it satisfies the following recurrence formula.
@@
(It has an explicit expression as well. See Lemma 2.1 for details.) Then, we have for $\mu$-a.e. $t$
@@
$\mathrm{dim}_{\mathrm{H}}\left((K_{1}+t)\cap K_{2}\right)=\frac{\lambda}{\log{7}}=0.356207\ldots\quad\left(\text{$\mu$-a.e. $t$}\right)$
@@
Informally, Kernel-expansion decomposes the integrand into a constant plus a uniformly convergent series of functions each having zero $\nu$-average. Note that by $\|\psi_{n}(\phi)\|_{\infty}\leq 2{\|\phi\|_{\infty}}^{n}$ and $\|\phi\|_{\infty}<1$, the uniform convergence of the right hand side in the above equation is always ensured.
@@
Define an infinite matrix $T=(b_{k,n})_{k,n\in\mathbb{N}_{0}}$ by $b_{k,0}=0$ for all $k$, and for $n>0$, as the coefficients appearing in the following expansion.
@@
Then, the pair $(D_{1},D_{2})$ is Kernel-expandable, and
@@
Let $b\geq 7$ be a natural number and $\mu$ a product measure. Consider $\tau,u\in\{0,1,\ldots,b-1\}$ and let
@@
This theorem shows that when $b\geq 7$ and we forbid only a single digit for each $D_{j}$, we have completed the calculation of $\mathrm{dim}_{\mathrm{H}}\left((K_{1}+t)\cap K_{2}\right)$. That is, either the already-known method by Kenyon and Peres can be used, or our new technique is viable.
Let us see an example where Kernel expansion yields the dimension for a non-degenerate case.
@@
In this case, any $\mu$-stationary measure is continuous. However, Kernel expansion is possible with $\phi(x)=\frac{7x+4}{4x+16}$. The operator norm of $T$ in the above “Theorem” is less than $\frac{649}{723}$ for this $\phi$. Define $a=(a_{n})_{n=0}^{\infty}\in\ell^{\infty}(\mathbb{N}_{0})$ by
@@
$\lambda=\sum_{n=0}^{\infty}\left(T^{n}a\right)_{0}=1.6363797884\ldots.$
@@
$\mathrm{dim}_{\mathrm{H}}\left((K_{1}+t)\cap K_{2}\right)=\frac{\lambda}{\log{7}}=0.8409328607\cdots$
@@
So, in this case, none of the $f_{i}$ is linear, and we do not see the type of symmetry in matrices $A_{i}$ that is present in the previous Example 1.10. Any stationary measure is continuous; yet $(D_{1},D_{2})$ is Kernel-expandable with $\phi(x)=\frac{-0.3914x-0.055}{-0.0639x+1}$, and we have
@@ 3.1  Introducing the method
Let $\phi:[-1,1]\to(-1,1)$ be a Möbius transformation, and suppose we have the following expansion for every natural number $n$:
@@ 3.2  Neumann Admissibility Condition and Kernel-expansion
In this subsection, we consider an analytic condition for $\phi$ so that the assumptions in Proposition 3.1 are satisfied. We call it the Neumann Admissibility Condition (or NAC), and assuming it, Theorem 3.2 calculates the top Lyapunov exponent of products of random matrices explicitly.
Let us recall the setting. Let $b\geq 2$ be an integer, and $\mu$ a probability measure on $\mathbb{T}=\mathbb{R}/\mathbb{Z}$, invariant and ergodic for $T_{b}x=bx\mathrm{\hskip 5.0ptmod\hskip 3.0pt}1$. We consider $b$ number of non-zero $2\times 2$ matrices with non-negative entries, $A_{0},\ldots,A_{b-1}$. The associated Möbius IFS $\{f_{i}\}_{i=0}^{b-1}$ is defined with
@@
satisfying the Neumann Admissibility Condition. Then $\{A_{0},\ldots,A_{b-1}\}$ is Kernel-expandable, and we can define a linear operator $T:\ell^{\infty}(\mathbb{N}_{0})\to\ell^{\infty}(\mathbb{N}_{0})$ and $a\in\ell^{\infty}(\mathbb{N}_{0})$ such that we have the following convergent series. $\left(\text{Here, $(\cdot)_{0}$ is the $0$-th entry.}\right)$
@@
$b_{k,n}=\sum_{i=0}^{b-1}\,\mu(t_{1}=i)\,\left(\frac{p_{2}(i)}{q_{2}(i)}\right)^{n}\,\left\{\sum_{\ell=1}^{\min\{k,n\}}\binom{n}{\ell}\binom{k-1}{\ell-1}\left(1-\frac{p_{1}(i)\,q_{2}(i)}{p_{2}(i)\,q_{1}(i)}\right)^{\ell}\left(-\frac{q_{1}(i)}{q_{2}(i)}\right)^{k}\right\}$
@@
$(\phi\circ f_{i})^{n}={\begin{pmatrix}A&B\\ C&D\end{pmatrix}\begin{pmatrix}\alpha(i)&\beta(i)\\ \gamma(i)&\delta(i)\end{pmatrix}{\begin{pmatrix}A&B\\ C&D\end{pmatrix}}^{-1}}\big(\phi\big)$
@@
This proves equation (3.6).
@@
Under the setting of  Theorem 3.2, for each $\mu$-valid index $i$, there is $a^{(i)}=\left(a_{n}^{(i)}\right)_{n=0}^{\infty}\in\ell^{\infty}(\mathbb{N}_{0})$ satisfying the following equality with uniformly convergent series.
@@
More precisely, for each $i$, the vector $a^{(i)}$ is given by
@@
For $0\leq i\leq b-1$, the integrand in (1.3) is expressed using $f_{i}$ as
@@ 3.3  Kernel expandability through Neumann series method
Theorem 1.9 (Restated). Let $b\geq 7$ be a natural number and $\mu$ a product measure. Consider $\tau,u\in\{0,1,\ldots,b-1\}$ and let
@@
Take an integer $b\geq 7$. Let $\tau\in\{1,2,\ldots,b-2\}$, and
@@
We are left to check the case $7\leq b\leq 21$.For these $b$ and $1\leq\tau\leq b-2$, we looked for the values of $A,B,C$ satisfying NAC for each pair of $(D_{1},D_{2})$, thus completing the proof.
@@
The pair $(D_{1},D_{2})$ is degenerate if one of $D_{1}$, $D_{2}$ misses one of the extreme digits 0,b10,b-1.
@@ 3.4  Evaluation of Neumann series method
$r=\max\left\{\max_{0\leq i\leq b-1}\left\{\left|\frac{C\delta(i)-D\gamma(i)}{A\delta(i)-B\gamma(i)}\right|\right\},\hskip 3.0pt\left|\frac{C}{A}\right|,\hskip 3.0pt\rho_{4}\right\}$ (3.11)
@@
$\left\|Ty-T_{N}y\right\|_{\infty}\leq R_{N}$
@@
$M=O\!\left(\log{1/\varepsilon}\right),\quad N=O\!\left(\log{1/\varepsilon}\right).$
@@
$O\!\left(\sum_{k=1}^{N}\sum_{n=1}^{N}\min{(k,n)}\right)=O\!\left(N^{3}\right)$

### v2

@@ 1.1  Background
$K_{j}=\left\{\sum_{n=1}^{\infty}\frac{d_{n}}{b^{n}}\hskip 2.0pt\middle|\hskip 2.0ptd_{n}\in D_{j}\right\}\subset\mathbb{T}=\mathbb{R}/\mathbb{Z}.$
@@
Let $\mu$ be a probability measure on $\mathbb{T}=\mathbb{R}/\mathbb{Z}$, invariant and ergodic for $T_{b}x=bx\mathrm{\hskip 5.0ptmod\hskip 3.0pt}1$. Define the $2$ by $2$ matrices $A_{0},\ldots,A_{b-1}$ by
@@
Except when $A_{t_{n}}\cdots A_{t_{1}}$ is the zero matrix for $\mu$-a.e. $t$ for some $n$, in which case the dimension is 0.
@@
This $\lambda$ is called the (top) Lyapunov exponent of the product of random matrices $A_{0},\ldots,A_{b-1}$. The RHS of (1.1) may not always exist, but it exists and is constant for $\mu$-a.e. $t$ ([2]). Also, the original statement in [8, Theorem 1.2] is formulated solely for the Lebesgue measure. However, a more general statement in [8, Corollary 5.8] guarantees that this is true for any invariant and ergodic probability measure.
@@
A probability measure $\nu$ on [1,1][-1,1] is said to be $\bm{\mu}$-stationary if for every continuous function $g$,
$\int_{[-1,1]}g\,d\nu=\int_{[-1,1]}\left(\int_{\mathbb{T}}g\circ f_{t_{1}}(x)\,d\mu(t)\right)\,d\nu(x).$
@@
Let $\mu$ be the Lebesgue measure, and $\nu$ be any $\mu$-stationary measure. The matrices $A_{0}$ and $A_{3}$ have rank one, so $\nu$ is unique and discrete (i.e. it is a countable sum of Dirac measures). The Lyapunov exponent can be calculated as
$\lambda=\frac{1}{6}\log{\frac{2}{3}}+\sum_{k=0}^{\infty}\frac{1}{4^{k+1}}\log{\frac{(3\cdot 2^{k})!}{(2^{k+1})!}}=0.797435\cdots.$
This gives the dimension to be $0.575228\cdots$.
@@
The pair $(D_{1},D_{2})$ is degenerate if and only if there is a $\mu$-valid index $i$ such that $A_{i}$ has rank one.
When this happens, the image of the corresponding $f_{i}$ is a singleton, making the $\mu$-stationary measure unique and discrete. We can then calculate the dimension by tracking the trajectory of the said singleton under all $f_{i}$. The calculations by Kenyon and Peres (such as Example 1.4) fall under this category. See the recent paper by Fan and Vebitskiy [4] for a general formula under this assumption.
@@ 1.2  Main results with examples
Let $\mu=\left(0,\frac{1}{2},0,\frac{1}{2},0,0,0\right)^{\otimes\mathbb{N}}$. Note that $f_{1}$ and $f_{3}$ have distinct attracting fixed point, $1$ and 0 respectively, so the semigroup generated by $\{A_{i}\}_{i}$ does not preserve any finite set of proper subspaces of $\mathbb{R}^{2}$. This implies the uniqueness of $\mu$-stationary measure $\nu$. Moreover, $\nu$ is continuous by [1, Lemma 4.2], so the method by Kenyon and Peres does not work. Take the $\mu$-stationary measure $\nu$ and let $\xi_{n}=\int\phi^{n}d\nu$. Then, it satisfies the following recurrence formula.
@@
(It has an explicit expression as well. See Lemma 2.1 for details.) Then, we have
@@
$\mathrm{dim}_{\mathrm{H}}\left((K_{1}+t)\cap K_{2}\right)=\frac{\lambda}{\log{7}}=0.356207\cdots\quad\left(\text{$\mu$-a.e. $t$}\right).$
@@
Note that by $\|\psi_{n}(\phi)\|_{\infty}\leq 2{\|\phi\|_{\infty}}^{n}$ and $\|\phi\|_{\infty}<1$, the uniform convergence of the right hand side in the above equation is always ensured.
Informally, Kernel-expansion decomposes the integrand into a constant plus a uniformly convergent series of functions each having zero $\nu$-average. We remark that if $\{A_{i}\}_{i}$ is Kernel-expandable, the integral in equation (1.3) takes the same value for every $\mu$-stationary measure. In other words, if there are $\mu$-stationary measures that yield different values of this integral, then $\{A_{i}\}_{i}$ is not Kernel-expandable.
@@
Let $\mu$ be a probability measure on $\mathbb{T}=\mathbb{R}/\mathbb{Z}$, invariant and ergodic for $T_{b}x=bx\mathrm{\hskip 5.0ptmod\hskip 3.0pt}1$. Define an infinite matrix $T=(b_{k,n})_{k,n\in\mathbb{N}_{0}}$ by $b_{k,0}=0$ for all $k$, and for $n>0$, as the coefficients appearing in the following expansion (where we assume such expansion exists).
@@
Then, the pair $(D_{1},D_{2})$ is Kernel-expandable with respect to $\mu$, and
@@
Let $b\geq 7$ be a natural number and $\mu$ Lebesgue measure. Consider $\tau,u\in\{0,1,\ldots,b-1\}$ and let
@@
This theorem shows that when $b\geq 7$ and we forbid only a single digit for each $D_{j}$, we have completed the calculation of $\mathrm{dim}_{\mathrm{H}}\left((K_{1}+t)\cap K_{2}\right)$ for Lebesgue a.e. $t$. That is, either the already-known method by Kenyon and Peres can be used, or our new technique is viable.
Let us see an example where Kernel-expansion yields the dimension for a non-degenerate case.
@@
Since $f_{0}$ and $f_{6}$ have distinct attracting fixed points, the semigroup generated by $\{A_{i}\}_{i}$ does not preserve any finite set of proper subspaces of $\mathbb{R}^{2}$, implying the uniqueness of $\mu$-stationary measure $\nu$. Furthermore, $\nu$ is continuous by [1, Lemma 4.2]. However, Kernel expansion is possible with $\phi(x)=\frac{7x+4}{4x+16}$. The operator norm of $T$ in the above “Theorem” is less than $\frac{649}{723}$ for this $\phi$. Define $a=(a_{n})_{n=0}^{\infty}\in\ell^{\infty}(\mathbb{N}_{0})$ by
@@
$\lambda=\sum_{n=0}^{\infty}\left(T^{n}a\right)_{0}=1.6363797884\cdots.$
@@
$\mathrm{dim}_{\mathrm{H}}\left((K_{1}+t)\cap K_{2}\right)=\frac{\lambda}{\log{7}}=0.8409328607\cdots.$
@@
So, in this case, none of the $f_{i}$ is linear, and we do not see the type of symmetry in matrices $A_{i}$ that is present in the previous Example 1.10. A similar discussion to that in Example 1.10 shows that the $\mu$-stationary measure is unique and continuous; yet $(D_{1},D_{2})$ is Kernel-expandable with $\phi(x)=\frac{-0.3914x-0.055}{-0.0639x+1}$, and we have
@@ 3.1  Introducing the method
Let $\phi:[-1,1]\to(-1,1)$ be a Möbius transformation, and $\mu$ a probability measure on $\mathbb{T}=\mathbb{R}/\mathbb{Z}$, invariant and ergodic for $T_{b}x=bx\mathrm{\hskip 5.0ptmod\hskip 3.0pt}1$. Suppose we have the following expansion for every natural number $n$:
@@ 3.2  Neumann Admissibility Condition and Kernel-expansion
In this subsection, we consider an analytic condition for $\phi$ so that the assumptions in Proposition 3.1 are satisfied for Lebesgue measure $\mu$. We call it the Neumann Admissibility Condition (or NAC), and assuming it, Theorem 3.2 calculates the top Lyapunov exponent of products of random matrices explicitly.
Let us recall the setting. Let $b\geq 2$ be an integer, and $\mu$ be a probability measure on $\mathbb{T}=\mathbb{R}/\mathbb{Z}$, invariant and ergodic for $T_{b}x=bx\mathrm{\hskip 5.0ptmod\hskip 3.0pt}1$. We consider $b$ number of non-zero $2\times 2$ matrices with non-negative entries, $A_{0},\ldots,A_{b-1}$. The associated Möbius IFS $\{f_{i}\}_{i=0}^{b-1}$ is defined with
@@
satisfying the Neumann Admissibility Condition. Then $\{A_{0},\ldots,A_{b-1}\}$ is Kernel-expandable with respect to $\mu$, and we can define a linear operator $T:\ell^{\infty}(\mathbb{N}_{0})\to\ell^{\infty}(\mathbb{N}_{0})$ and $a\in\ell^{\infty}(\mathbb{N}_{0})$ such that we have the following convergent series. $\left(\text{Here, $(\cdot)_{0}$ is the $0$-th entry.}\right)$
@@
$b_{k,n}=\sum_{i=0}^{b-1}\,\mu(t_{1}=i)\,\left(\frac{p_{2}(i)}{q_{2}(i)}\right)^{n}\,\left(-\frac{q_{1}(i)}{q_{2}(i)}\right)^{k}\,\left\{\sum_{\ell=1}^{\min\{k,n\}}\binom{n}{\ell}\binom{k-1}{\ell-1}\left(1-\frac{p_{1}(i)\,q_{2}(i)}{p_{2}(i)\,q_{1}(i)}\right)^{\ell}\right\}$
@@
$(\phi\circ f_{i})^{n}=\left({\begin{pmatrix}A&B\\ C&D\end{pmatrix}\begin{pmatrix}\alpha(i)&\beta(i)\\ \gamma(i)&\delta(i)\end{pmatrix}{\begin{pmatrix}A&B\\ C&D\end{pmatrix}}^{-1}}\big(\phi\big)\right)^{n}$
@@
The convergence is assured by NAC. This proves equation (3.6).
@@
Under the setting of  Theorem 3.2, for each $\mu$-valid $0\leq i\leq b-1$, there is $a^{(i)}=\left(a_{n}^{(i)}\right)_{n=0}^{\infty}\in\ell^{\infty}(\mathbb{N}_{0})$ satisfying the following equality with uniformly convergent series.
@@
More precisely, for each $\mu$-valid $0\leq i\leq b-1$, the vector $a^{(i)}$ is given by
@@
For $\mu$-valid $0\leq i\leq b-1$, the integrand in (1.3) is expressed using $f_{i}$ as
@@ 3.3  Kernel expandability through Neumann series method
Theorem 1.9 (Restated). Let $b\geq 7$ be a natural number and $\mu$ Lebesgue measure. Consider $\tau,u\in\{0,1,\ldots,b-1\}$ and let
@@
Take an integer $b\geq 7$, and let $\mu$ be Lebesgue measure. Let $\tau\in\{1,2,\ldots,b-2\}$, and
@@
We are left to check the case $7\leq b\leq 21$.For these $b$ and $1\leq\tau\leq b-2$, we rigorously proven the existence of the values of $A,B,C$ satisfying NAC for each pair of $(D_{1},D_{2})$, thus completing the proof.
@@
Let $\mu$ be Lebesgue measure. The pair $(D_{1},D_{2})$ is degenerate if one of $D_{1}$, $D_{2}$ misses one of the extreme digits 0,b10,b-1.
@@ 3.4  Evaluation of Neumann series method
$r=\max\left\{\max_{\begin{subarray}{c}0\leq i\leq b-1\\ i:\,\mu\text{-valid}\end{subarray}}\left\{\left|\frac{C\delta(i)-D\gamma(i)}{A\delta(i)-B\gamma(i)}\right|\right\},\hskip 3.0pt\left|\frac{C}{A}\right|,\hskip 3.0pt\rho_{4}\right\}.$ (3.11)
@@
$\left\|Ty-T_{N}y\right\|_{\infty}\leq R_{N}.$
@@
$M=O\!\left(\log{(1/\varepsilon)}\right),\quad N=O\!\left(\log{(1/\varepsilon)}\right).$
@@
$O\!\left(\sum_{k=1}^{N}\sum_{n=1}^{N}\min{(k,n)}\right)=O\!\left(N^{3}\right).$
@@ Acknowledgement
I had help from ChatGPT (OpenAI) to develop the codes used in the proof. All codes were reviewed and integrated by me, who is responsible for the final implementation and results.